# A Proactive Decoy Selection Scheme for Cyber Deception using MITRE ATT&CK

### v1

@@ 4.1 Threat Model
    $A^{AND}$ is the subset of techniques with type $AND$ preconditions.
@@
    $A^{OR}$ is the subset of techniques with type $OR$ preconditions.
@@
    If $v_{i},i\in A$ and $v_{j},j\in A$ the edge $(v_{i},v_{j})$ indicates that technique $i$ is a precondition for the execution of technique $j$. For example, the technique Hijack Execution Flow, which enables to run malicious code by manipulating how the operating system run programs, must be first executed by the attacker in order to employ the technique OS Credential Dumping, which allows locating and stealing credential material from the operating system and software.
@@
The resulting attack graph can be considered as an augmented version of the original MITRE ATT&CK Enterprise matrix that models the dependency between each technique in terms of execution preconditions as well as their impact on the target environment. In detail, we can assess how an attacker can practically achieve the execution of a technique in the environment by inspecting the predecessor nodes in the graph. The same reasoning applies when analyzing how an outcome can be achieved starting from a technique.
@@
Figure 1: Example of attack path when nodes of type AND are included. Purple arrows represent OR preconditions. Orange arrows represent AND preconditions. Red dashed arrows indicate the attack path from node $v_{s}$ to node $v_{t}$.
@@ 4.2 Defender Model
To implement a proactive cyber deception approach, the defender leverages the proposed attack graph to model the expected adversary behavior by means of attack paths, that indeed express which ATT&CK techniques are required and how they are correlated with one another to accomplish some goal. In detail, we assume that the defender is unaware of the attacker goals, hence he selects a set of technique outcomes $T\subseteq O$ that would severely damage the system once accomplished and considers them as attack targets of a hypothetical adversary. Based on this selection, the defender analyzes which attack steps are required to reach the various targets from a set of techniques and/or outcomes $S=\{s:s\in A\cup O\land s\notin T\}$, that model the attacker initial capabilities. The selection of the elements in $S$ can follow different criteria. For example, the defender could be interested in tracking the adversary progression towards his goals starting from different attack stages. Using as reference Fig. 1, the node $v_{s}$ could correspond to the technique Exploitation of Remote Services that indicates the capability of an attacker to move laterally between internal system components of the defender environment exploiting his production services. This scenario indicates an attacker that has already gained persistence within the system by evading security countermeasures. The attack paths originating from such node are likely to employ more advanced attack tactics such as Command and Control and Exfiltration. Differently, the analysis of attack paths including early attack stages could be indicated by configuring node $v_{s}$ with the technique Shared Module. The latter indicates an adversary expanding his initial foothold within the system by exploiting shared executable files between different process to run arbitrary code.
@@
The defender employs the resulting threat profile to optimize the selection of decoys to deploy within its system. In particular, we assume that the defender has the availability to deploy decoys that are specifically tailored to expose an attacker using each technique in the threat profile. In other words, for every ATT&CK technique, there exists a decoy suited for its detection. Our assumption is supported by MITRE frameworks, namely Engage [28] and Def3nd [29], that offer tools and resources to increase the effectiveness of adversary engagement strategies like cyber deception. In particular, they provide a mapping between ATT&CK techniques and possible decoys in order to facilitate the design of cyber deception solutions for cybersecurity practitioners. Although this mapping is not exhaustive for all ATT&CK techniques, we consider that it exists for any technique to simplify the design of the proposed scheme. This is still a reasonable assumption as any software artifact that can mimic a functionality of the legitimate system can serve as decoy. Consequently, any ATT&CK technique that has been observed to target or leverage those system components can be exposed by a suitably crafted decoy. Nonetheless, we acknowledge that this reasoning neglects the actual implementation feasibility required to create such decoys that might be prohibitive for real-world scenarios. For this reason, we leave as a future analysis the investigation of a decoy selection solution where only a limited subset of the adversary techniques can be exposed by a decoy.
@@
    Attack surface minimization: the number of allocated decoys is minimized to limit the attack surface increase. The introduction of any new component (either software or hardware) in an IT infrastructure can be potentially exploited by an attacker to further compromise the environment. Therefore, it is a good cybersecurity practice to generally restrict the injection of new elements to limit the attack surface increase [30]. This reasoning still applies to decoys that may be possibly intertwined with the production environment and thus could be exploited unpredictably to damage legitimate system components.
@@ 5.1 Optimal decoy selection
We analytically compute the selection of deceptive artifact by proposing an integer optimization problem that can be considered as a variant of the Multi-Terminal Vertex Separator problem [31]. The general idea is to partition the graph vertices into three different subsets $X,Y$ and $Z$ such that vertices in $X$, if removed, disconnect all paths from any vertex in $Y$ to any vertex in $Z$. We indicate the node assignment to a partition by defining the binary indicator variables $x_{i},y_{i},z_{i}$ that individually assume value $1$ if vertex $v_{i}\in X$, $v_{i}\in Y$, $v_{i}\in Z$, respectively, and value 0 otherwise.
The set $Y$ and $Z$ represent the source and target partitions and are initialized by setting $y_{i}=1,\forall v_{i}\in S$ and $z_{i}=1,\forall v_{i}\in T$, respectively. Differently, the $X$ is initialized as an empty set and represents the decoys exposing techniques disconnecting $Y$ from $Z$. Every vertex must be assigned to exactly one partition, hence all the unnecessary vertices not belonging to $X$ are either assigned to $Y$ or $Z$. To correctly compute the partition $X$, we need to account for the type of every node. Nodes of type AND impose the reachability of all their predecessors in order to be employed in the attack paths, whereas nodes of type OR require the reachability of at least one of their predecessors. We encode this information by further defining the binary indicator variable $r_{i,s}$ that assumes value 1 if node $v_{i}\in R_{\bar{G}}(v_{s}),\forall v_{s}\in S$ (i.e., if node $v_{i}$ is reachable from any node $v_{s}\in S$). Finally, we introduce the variable $c_{i}$ that assume a fixed value of $1$ if $i\in A^{\bar{M}}$ and variable value $\beta\geq 1$ if $i\in A^{M}$. This variable allows associating a higher cost to techniques with mitigations in order to disincentivize their selection during the computation of the partition set $X$. Leveraging the introduced variables, the optimization problem can be written as
@@
Constraint (3) and (4) implement the reachability logic condition for each vertex based on its type. In detail, they ensure that each node is labelled as reachable from a source node in $S$, i.e. $r_{i,s}=1$, only if its predecessors can be reached as well. More precisely, (3) guarantees that a OR vertex $v_{i}$ is reachable if it is assigned to the partition $Y$ or it is assigned to the $Z$ partition and at least one of its predecessors, indicated by the set $N_{\bar{G}}(v_{i})$, is reachable. Note that if $v_{i}$ is assigned to the partition $X$ is unreachable by definition, hence $r_{i,s}=0$ due to $y_{i}+z_{i}=0$. Instead, (4) guarantees the reachability of a AND vertex by assessing the opposite condition, i.e. when a node of type AND is not reachable (we considered the unreachable condition for the nodes of type AND as it allow us to employ the same variables without the need of introducing new ones). This condition occurs when a node of type AND belongs to $X$ partition and has at least one unreachable predecessor. Constraint (5) enforces that the nodes in the $X$ partition must disconnect any attack path between each $v_{i}\in Y$ and $v_{j}\in Z$ given the reachability of the various nodes. In particular, it forbids the existence of a directed edge $(v_{i},v_{j})\in\bar{E}$ such that $v_{i}\in Y$ and $v_{j}\in Z$ since this would violate the property of the $X$ partition unless either $v_{i}$ or $v_{j}$ is unreachable. Constraint (6) initializes the $Y$ and $Z$ subsets with the vertexes in $S$ and $T$, respectively. Moreover, it ensures that subset $X$ does not contain any technique outcome as they are not eligible to be detected by a decoy. Constraint (7) guarantees that every vertex in the graph is assigned to one partition only. Finally, constraint (8) enforces the integer nature of the problem formulation.
The above formulation (2)-(8) has complexity NP-Hard as it can be reduced to the original Multi-Terminal Vertex Separator problem in polynomial time when node types are ignored. Note that the non-linearity introduced by the min operator in constraints (3)-(4) can be linearized by using suitable auxiliary variables as explained in [32]. In general, the non-polynomial complexity makes the solution computation unpractical for large graphs due to the high execution time. However, the number of nodes composing $\bar{G}$ is limited since it is constrained by the maximum number of available ATT&CK techniques, hence setting an upper bound to the scaling complexity. Therefore, as we are going to show in the performance evaluation section, the design of a heuristic approximating the optimal solution is unnecessary due to the negligible resolution time that we experimentally observed during the simulations.
Figure 2: Example of decoy selection solution. The graph represents the threat profile $\bar{G}$ extracted by the original attack graph $G$. The green vertex represents the attacker initial capabilities and indicates the source node of the attack paths, the red vertex represents the attacker target. Each yellow vertex represents a decoy exposing the indicated technique.
@@ 5.2 Decoy selection solution overview
We discuss an example of decoy selection solution to better showcase the various steps on a practical scenario and to provide a better understanding of the proposed methodology. We provide a visual representation of the solution in Fig. 2. We consider a defender that wants to allocate a number of decoys to expose a malicious user, that has already a foothold within the environment and that aims at executing malicious code to infect a host within the network. The deployed decoys are selected according to the procedure described in the previous section. In detail, we identify the attacker target as the set $T=\{infectedComputer\}$, where infectedComputer is a technique outcome indicating an attacker that has successfully run malicious software in a host. Similarly, we model the attacker initial foothold by initializing the source node set as $S=\{userRights\}$, where userRights is a technique outcome indicating an attacker that has acquired user right privileges and it is capable of interacting with most of the system assets. Given the defined $S$ and $T$, we can compute the threat profile $\bar{G}$ by first computing all the attack paths with (1) and then pruning the unnecessary nodes and related edges from $G$. Finally, we compute the minimum number of nodes that disconnect $S$ from $T$ leveraging (2)-(8). Each node corresponds to a decoy that detects the corresponding technique. In the considered example, the solution corresponds to decoys exposing shortcutModification and rightToLeftOverride techniques. The former indicates an adversary modifying the shortcuts within the startup folder so that malicious code is executed upon system boot. A possible deceptive artifact may be a fake startup folder that treats any modification of its content as a malicious activity. The latter consists in the usage of the “right-to-left override” Unicode character to trick a legitimate user into opening malicious files. A possible deceptive artifact may be an agent that monitors files containing such character and that redirects their execution within a controlled and isolated environment to waste attacker resources in fake interactions. The resulting decoy selection solution takes advantage of the node types to minimize the number of decoys. As a matter of fact, the selection of a decoy to detect a precondition enabling the maliciousFile attack step can efficiently detect any attack path. This is due to the fact that maliciousFile is a node of type AND, hence it requires all its predecessors to be reachable from the source node in order to be employable by an adversary (in the cases of maliciousFile was a node of type OR, multiple decoys would be required, one for each path reaching its predecessors). We remark that this is an example meant to highlight the features of the proposed scheme in a simple scenario. The number of nodes composing the threat profile $\bar{G}$ largely depends on the number of considered source nodes and attack targets. More complex configurations generate threat profiles potentially composed by hundreds of nodes that offer less trivial decoy selection solutions and thus justify the need to provide an analytical problem formulation.
@@ 6.1 Simulation Setup
The simulation environment was developed in Python. In particular, we used NetworkX library to handle all the basic graph operations, whereas we implemented custom methods to handle the computation involving nodes of type AND. We solved (2)-(8) using the SCIP optimization suite, an open-source solver for integer linear programming [33]. We assessed the performance of the optimal decoy selection, referred to as Optimal, with the following baseline schemes:
@@
We compare the security performance of the proposed schemes using three main metrics that captures a different feature of the resulting decoy selection:
@@
    Attack path interception ratio: this metric measures the deception performance of the selected decoys in luring attackers before they reach their goal once deployed. We compute its value as the ratio between the number of attack paths containing at least one technique that can be exposed by a decoy and the total number of attack paths in $\bar{G}$. A high value indicates an effective decoy placement that is tailored to the attacker behavior modelled using ATTA&CK techniques.
@@
We considered different configuration of attack targets ranging from 1 to 9 and we averaged the results of 100 random instances of threat profiles $\bar{G}$ generated as follows. We modelled the initial attacker capabilities assuming $S=\{userRight\}$ as this reflects a realistic scenario where an attacker has gained user privileges within the system and leverages those privileges to further penetrate the environment with more advanced techniques. Although this is not comprehensive of every possible attack path, the majority of real-world attacks reported in the MITRE ATT&CK framework shares this intermediate attack goal. As a matter of fact, a malicious user with no level of rights has a limited pool of available techniques which are indeed focused to gain some permission rights within the system, thus allowing to expand his capabilities. We selected the various attack targets by randomly sampling a number of technique outcomes from the set $O$ for each configuration. Practically, we extracted the initial set of technique outcomes by individually inspecting each attack step encoded in enterpriseLang to check if it was composed by additional elements other than the technique name. This procedure produced 153 technique outcomes which were considered eligible to be selected as attack targets.

### v2

@@ 4.1 Threat Model
    $A^{AND}$ is the subset of techniques with preconditions of type $AND$.
@@
    $A^{OR}$ is the subset of techniques with preconditions of type $OR$.
@@
    If $v_{i},i\in A$ and $v_{j},j\in A$ the edge $(v_{i},v_{j})$ indicates that technique $i$ is a precondition for the execution of technique $j$. For example, the technique Hijack Execution Flow, which enables to run malicious code by manipulating how the operating system runs programs, must be first executed by the attacker to employ the technique OS Credential Dumping, which allows the location and theft of credentials material from the operating system and software.
@@
The resulting attack graph can be considered as an augmented version of the original MITRE ATT&CK Enterprise matrix that models the dependency between each technique in terms of execution preconditions as well as their impact on the target environment. In detail, we can assess how an attacker can practically achieve the execution of a technique in the environment by inspecting the predecessor nodes in the graph. The same reasoning applies when analyzing how an outcome can be achieved starting from a technique. In Fig. 1, we show a pictorial overview of how we combine the information provided by the MITRE ATT&CK framework and enterpriseLang to build the proposed threat model.
@@
Figure 2: Example of attack path when AND nodes are included. Purple arrows represent OR preconditions. Orange arrows represent AND preconditions. Red dashed arrows indicate the attack path from node $v_{s}$ to node $v_{t}$.
@@ 4.2 Defender Model
To implement a proactive cyber deception approach, the defender leverages the proposed attack graph to model the expected adversary behavior by means of attack paths, that indeed express which ATT&CK techniques are required and how they are correlated with one another to accomplish some goal. In detail, we assume that the defender is unaware of the attacker goals, hence he selects a set of technique outcomes $T\subseteq O$ that would severely damage the system once accomplished and considers them as attack targets of a hypothetical adversary. Based on this selection, the defender analyzes which attack steps are required to reach the various targets from a set of techniques and/or outcomes $S=\{s:s\in A\cup O\land s\notin T\}$, that model the attacker initial capabilities. The selection of the elements in $S$ can follow different criteria. For example, the defender could be interested in tracking the adversary progression towards his goals starting from different attack stages. Using as reference Figure 2, node $v_{s}$ could correspond to the technique Exploitation of Remote Services that indicates the capability of an attacker to move laterally between internal system components of the defender environment exploiting his production services. This scenario indicates an attacker that has already gained persistence within the system by evading security countermeasures. The attack paths originating from such node are likely to employ more advanced attack tactics such as Command and Control and Exfiltration. Differently, the analysis of attack paths including early attack stages could be indicated by configuring node $v_{s}$ with the technique Shared Module. The latter indicates an adversary expanding his initial foothold within the system by exploiting shared executable files between different process to run arbitrary code.
@@
The defender employs the resulting threat profile to optimize the selection of decoys to deploy within its system. In particular, we assume that the defender has the availability to deploy decoys that are specifically tailored to expose an attacker using each technique in the threat profile. In other words, for every ATT&CK technique, there exists a decoy suited for its detection. Our assumption is supported by MITRE frameworks, namely Engage (MITRE Engage, 2022) and Def3nd (Kaloroumakis and Smith, 2021), that offer tools and resources to increase the effectiveness of adversary engagement strategies like cyber deception. In particular, they provide a mapping between ATT&CK techniques and possible decoys in order to facilitate the design of cyber deception solutions for cybersecurity practitioners. Although this mapping is not exhaustive for all ATT&CK techniques, we consider that it exists for any technique to simplify the design of the proposed scheme. This is still a reasonable assumption as any software artifact that can mimic a functionality of the legitimate system can serve as decoy. Consequently, any ATT&CK technique that has been observed to target or leverage those system components can be exposed by a suitably crafted decoy. Nonetheless, we acknowledge that this reasoning neglects the actual implementation feasibility required to create such decoys that might be prohibitive for real-world scenarios. For this reason, we leave as a future analysis the investigation of a decoy selection solution where only a limited subset of the adversary techniques can be exposed by a decoy.
@@
    Attack surface minimization: the number of allocated decoys is minimized to limit the attack surface increase. The introduction of any new component (either software or hardware) in an IT infrastructure can be potentially exploited by an attacker to further compromise the environment. Therefore, it is a good cybersecurity practice to generally restrict the injection of new elements to limit the attack surface increase (Theisen et al., 2018). This reasoning still applies to decoys that may be possibly intertwined with the production environment and thus could be exploited unpredictably to damage legitimate system components.
@@ 5 Problem formulation
Figure 3: Example of decoy selection solutions when AND nodes and OR nodes are both considered (AND+OR solution, yellow nodes) and when OR nodes only are considered (OR solution, purple nodes).
@@ 5.1 Optimal decoy selection
We analytically compute the selection of deceptive artifacts by proposing an integer optimization problem that can be considered as a variant of the Multi-Terminal Vertex Separator problem (Cornaz et al., 2019). The general idea is to partition the graph vertices into three different subsets $X,Y$ and $Z$ such that vertices in $X$, if removed, disconnect all paths from any vertex in $Y$ to any vertex in $Z$. We indicate the node assignment to a partition by defining the binary indicator variables $x_{i},y_{i},z_{i}$ that individually assume value $1$ if vertex $v_{i}\in X$, $v_{i}\in Y$, $v_{i}\in Z$, respectively, and value 0 otherwise.
The set $Y$ and $Z$ represent the source and target partitions and are initialized by setting $y_{i}=1,\forall v_{i}\in S$ and $z_{i}=1,\forall v_{i}\in T$, respectively. Differently, partition $X$ is initialized as an empty set and represents the decoys exposing techniques disconnecting $Y$ from $Z$. Every vertex must be assigned to exactly one partition, hence all the unnecessary vertices not belonging to $X$ are either assigned to $Y$ or $Z$. To correctly compute the partition $X$, we need to account for the type of every node. Nodes of type AND impose the reachability of all their predecessors in order to be employed in the attack paths, whereas nodes of type OR require the reachability of at least one of their predecessors. We encode this information by further defining the binary indicator variable $r_{i,s}$ that assumes value 1 if node $v_{i}\in R_{\bar{G}}(v_{s}),\forall v_{s}\in S$ (i.e., if node $v_{i}$ is reachable from any node $v_{s}\in S$). Finally, we introduce the variable $c_{i}$ that assume a fixed value of $1$ if $i\in A^{\bar{M}}$ and variable value $\beta\geq 1$ if $i\in A^{M}$. This variable allows associating a higher cost to techniques with mitigations in order to disincentivize their selection during the computation of the partition $X$. Leveraging the introduced variables, the optimization problem can be written as
@@
Constraint (3) and (4) implement the reachability logic condition for each vertex based on its type. In detail, they ensure that each node is labeled as reachable from a source node in $S$, i.e. $r_{i,s}=1$, only if its predecessors can be reached as well. More precisely, (3) guarantees that a OR vertex $v_{i}$ is reachable if it is assigned to the partition $Y$ or it is assigned to the $Z$ partition and at least one of its predecessors, indicated by the set $N_{\bar{G}}(v_{i})$, is reachable. Note that if $v_{i}$ is assigned to the partition $X$, it is unreachable by definition, hence $r_{i,s}=0$ due to $y_{i}+z_{i}=0$. Instead, (4) guarantees the reachability of a AND vertex by assessing the opposite condition, i.e. when it is not reachable (we considered the unreachable condition for the nodes of type AND as it allow us to employ the same variables without the need of introducing new ones). This condition occurs when a node of type AND belongs to $X$ partition and has at least one unreachable predecessor. Constraint (5) enforces that the nodes in the $X$ partition must disconnect any attack path between each $v_{i}\in Y$ and $v_{j}\in Z$ given the reachability of the various nodes. In particular, it forbids the existence of a directed edge $(v_{i},v_{j})\in\bar{E}$ such that $v_{i}\in Y$ and $v_{j}\in Z$ since this would violate the property of the $X$ partition unless either $v_{i}$ or $v_{j}$ is unreachable. Constraint (6) initializes $Y$ and $Z$ with the vertexes in $S$ and $T$, respectively. Moreover, it ensures that $X$ does not contain any technique outcomes as they are not eligible to be selected as decoys. Constraint (7) guarantees that every vertex in the graph is assigned to one partition only. Finally, constraint (8) enforces the integer nature of the problem formulation.
The above formulation (2)-(8) has complexity NP-Hard as it can be reduced to the original Multi-Terminal Vertex Separator problem in polynomial time when nodes type is ignored. Note that the non-linearity introduced by the min operator in constraints (3)-(4) can be linearized by using suitable auxiliary variables as explained in (Belov et al., 2016). In general, the non-polynomial complexity makes the solution computation unpractical for large graphs due to the high execution time. However, the number of nodes composing $\bar{G}$ is limited since it is constrained by the maximum number of available ATT&CK techniques, hence setting an upper bound to the scaling complexity. Therefore, as we are going to show in the performance evaluation section, the design of a heuristic approximating the optimal solution is unnecessary due to the negligible resolution time that we experimentally observed during the simulations.
@@ 5.2 Decoy selection solution overview
We discuss an example of decoy selection solution to better showcase the various steps on a practical scenario and to provide a better understanding of the proposed methodology. We provide a visual representation of the solution in Fig. 3. We consider a defender that wants to allocate a number of decoys to expose a malicious user that has already a foothold within the environment and that aims at executing malicious code to infect a host within the network, as well as to steal user credentials. The deployed decoys are selected according to the procedure described in the previous section. In detail, we identify the attacker target as the set $T=\{infectedComputer,userCredential\}$, where infectedComputer and userCredential are technique outcomes indicating an attacker that has successfully run malicious software on a host and that has managed to steal account name and passwords of a legitimate user. Similarly, we model the attacker initial foothold by initializing the source node set as $S=\{userRights\}$, where userRights is a technique outcome indicating an attacker that has acquired user right privileges and he is capable of interacting with most of the system assets. Given the defined $S$ and $T$, we can compute the threat profile $\bar{G}$ by first computing all the attack paths with (1) and pruning the unnecessary nodes and related edges from $G$. Then, we compute the minimum number of nodes that disconnect $S$ from $T$ solving (2)-(8) under two different scenarios in order to better highlight the benefit of considering execution preconditions of type AND to reduce the number of decoys. Specifically, we compare the solution obtained when AND nodes are correctly considered with respect to a solution computed assuming that each node in $\bar{G}$ is of type OR only. Note that a solution which assumes each node to be of type AND would not ensure the interception of all attack paths. For example, a decoy exposing any of the execution preconditions needed to achieve an attack target of type OR would be wrongly considered as optimal. For this reason we ignore this scenario as it would provide sub-optimal solutions that do not add any new meaningful insight to our approach. We visually represent the decoys selected by the two solutions in Fig. 3 using yellow nodes for the “AND+OR solution” and purple nodes for “OR solution”, respectively. Note that double-colored nodes indicate that both solutions share the same decoy selection.
When AND and OR nodes are both considered, the optimal solution corresponds to three decoys exposing shortcutModification, rightToLeftOverride and exploitationForPrivilegeExalation techniques. In detail, shortcutModification indicates an adversary modifying the shortcuts within the startup folder so that malicious code is executed upon system boot. A possible deceptive artifact may be a fake startup folder that treats any modification of its content as a malicious activity. The technique rightToLeftOverride consists in the usage of the “right-to-left override” Unicode character to trick a legitimate user into opening malicious files. A possible deceptive artifact may be an agent that monitors files containing such character and that redirects their execution within a controlled and isolated environment to waste attacker resources in fake interactions. Finally, exploitationForPrivilegeExalation denotes the utilization of any software vulnerabilities and/or system misconfiguration to gain administrator privileges. A possible deceptive artifact can consist of the deployment of an application with vulnerabilities that have been exploited to escalate privileges in order to lure and trap the attacker within an isolated environment. In contrast, when every node is assumed to be of type OR, the optimal solution selects five decoys (two more than the previous case) exposing the techniques exploitationForPrivilegeExalation, indicatorRemoval, shortcutModification, spearphisingAttachment and maliciousFile.
The difference between the two solutions is explained by the role that AND nodes play in the solution computation. Intuitively, AND nodes are used by the optimal formulation to cleverly select OR nodes which can effectively minimize the number of required decoys. In the considered example, the selection of exploitationForPrivilegeExalation and rightToLeftOverride techniques computed by the “AND+OR solution” ensures the interception of any attack path containing clearEventLogs and maliciousFile techniques. As a matter of fact, the latter require AND preconditions, hence a proper choice of their predecessors can simultaneously intercept the related attack paths as well as any other attack paths composed by OR nodes. Conversely, the “OR solution”, which ignores AND preconditions, must select multiple predecessors of AND nodes in order to prevent any possible attack paths, increasing the number of required decoys. This behavior is further exacerbated in more complex threat profiles composed of multiple attack targets, hence an effective decoy selection exploiting the property of AND nodes can greatly reduce the number of needed decoys. Moreover, we further remark that this is an example meant to highlight the features of the proposed scheme in a simple scenario. The number of nodes composing the threat profile $\bar{G}$ largely depends on the number of considered source nodes and attack targets. More complex configurations generate threat profiles potentially composed by hundreds of nodes that offer less trivial decoy selection solutions and thus justify the need to provide an analytical problem formulation.
@@ 6.1 Simulation Setup
The simulation environment was developed in Python. In particular, we used NetworkX library to handle all the basic graph operations, whereas we implemented custom methods to handle the computation involving nodes of type AND. We solved (2)-(8) using the SCIP optimization suite, an open-source solver for integer linear programming (Bestuzheva et al., 2023). We assessed the performance of the optimal decoy selection, referred to as Optimal, with the following baseline schemes:
@@
We compare the security performance of the proposed schemes using three main metrics that capture a different feature of the resulting decoy selection:
@@
    Attack path interception ratio: this metric measures the deception performance of the selected decoys in luring attackers before they reach their goal once deployed. We compute its value as the ratio between the number of attack paths containing at least one technique that can be exposed by a decoy and the total number of attack paths in $\bar{G}$. A high value indicates an effective decoy placement that is tailored to the attacker behavior modeled using ATTA&CK techniques.
@@
We considered different configurations of attack targets ranging from 1 to 9 and averaged the results of 100 random instances of threat profiles $\bar{G}$ generated as follows. We modeled the initial attacker capabilities assuming $S=\{userRight\}$ as this reflects a realistic scenario where an attacker has gained user privileges within the system and leverages those privileges to further penetrate the environment with more advanced techniques. Although this is not comprehensive of every possible attack path, the majority of real-world attacks reported in the MITRE ATT&CK framework shares this intermediate attack goal. As a matter of fact, a malicious user with no level of rights has a limited pool of available techniques which are indeed focused to gain some permission rights within the system, thus allowing to expand his capabilities. We selected the various attack targets by randomly sampling a number of technique outcomes from the set $O$ for each configuration. Practically, we extracted the initial set of technique outcomes by individually inspecting each attack step encoded in enterpriseLang to check if it was composed by additional elements other than the technique name. This procedure produced 153 technique outcomes which were considered eligible to be selected as attack targets.